OpenStudy (anonymous):

If c does not equal 0 and the slope of the line passing through (-c, 2c) and (4c, a) is 1, what is the expression for a in terms of c?

OpenStudy (anonymous):

slope m between two points (x1,y1) (x2,y2) is given by : m = (y2-y1) / (x2-x1) in our case (x1,y1) = (-c,2c) (x2,y2) = (4c,a) m = 1 so : 1 = (a-2c) / (4c -(-c)) 1 = (a-2c) / (5c) 5c = a - 2c a = 7c

OpenStudy (anonymous):

thank you :)